OpenStudy (anonymous):

TWO CYLINDERS ARE SIMILAR. Radio of surface areas is 25:36 ... what is the ratio of the radius of the smaller cylinder to the radius of the larger cylinder? what is the ratio of the volume of smaller cylinder to the volume of the larger cylinder?

OpenStudy (anonymous):

Ratio of corresponding parts: \(\Large\frac{a}{b}\) Ratio of surface areas: \(\Large\frac{a^2}{b^2}\) Ratio of volumes: \(\Large\frac{a^3}{b^3}\)